OpenStudy (anonymous):

i need help ;/ if x equals -6, then find the matching value for y. -x+y=-3

OpenStudy (anonymous):

You replace the value of x in your equation -(-6)+y=-3 6+y=-3 y=-9

OpenStudy (anonymous):

-(-6)+y=-3 y=-6-3 y=-9

OpenStudy (anonymous):

thanks ;]